Text Size:
Smaller Text Normal Text Larger Text

Facebook page

Twitter profile

YouTube channel


P69-502: Planets around Subgiants

P.I.: Marcus Lovmand Marcussen

This is a Monitoring program using FIES.

Instrument setup

  • FIES

  • High-res fiber F4

Afternoon instructions

  • Run afternoon calibrations: fies-calibs -fib 4

Observing instructions

  • On the Queue page, make sure the OB-generated observing script is written to disk, or else 'compile'

  • Go to directory: cd ~/obs/scripts/69-502/ and do ls -lrt

  • Start the requested OB generated script

Program Tracking
Proposal 69-502
Semester P69: 20240401 - 20241001
Principal Investigator Marcus Lovmand Marcussen
Programme Planets around Subgiants
Instrument FIES
Time Allocation 15 sets of obs., 1.5h/obs.
Total Nights 1.5
Type Mon
Notes Fixed dates to be scheduled
Total Time Spent 08:18 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
69-502 2024-04-02 01:10 Mon 24sex, hd82886, hd99706, hd108863,
hd131496, hd192699, hd200964
FIHd020094 - FIHd020110, FIHd020156 - FIHd020161 All targets except for hd33844 were observed. hd33844 was not observed due to high humidity. Clear skies, observed with normal conditions Observed
69-502 2024-04-04 01:08 Mon hd82886, 24sex, hd99706, hd108863,
hd131496, hd192699, hd200964
FIHd040074 - FIHd040091, FIHd040096 - FIHd040099, FIHd040104 - FIHd040106 Observed instead of proposal 69-010 targets as requested due to issues in the planned scripts. All the targets were observed during a clear sky and seeing 1.0''-2.00'' Observed
69-502 2024-04-13 01:04 Mon 24sex, hd82886, hd99706, hd108863,
hd131496, hd200964, hd192699
FIHd130094 - FIHd130109, FIHd130136 - FIHd130141 All the targets were observed successfully. Some of the OBs had to be restarted due to stalling, which happened to some other FIES OBs as well tonight. The weather conditions were dusty, but cloudless. Observed
69-502 2024-04-25 00:53 Mon 24sex, hd82886, hd99706, hd108863,
FIHd250063 - FIHd250079 Thin clouds, and good seeing. We were not able to observe the last two targets before sunrise Observed
69-502 2024-05-04 01:08 Mon 24sex, hd82886, hd99706, hd108863,
hd131496, hd192699, hd200964
FIHe040094 - FIHe040105, FIHe040112 - FIHe040114, FIHe040125 - FIHe040130 Observed in thin clouds with seeing below 1.0" Observed
69-502 2024-05-13 00:20 Mon hd200964, hd192699 FIHe130089 - FIHe130094 Targets observed with clear skies, 2" seeing. Be cautious with calibration frames Observed
69-502 2024-05-19 01:12 Mon 24sex, hd82886, hd99706, hd108863,
hd131496, hd192699, hd200964
FIHe190094 - FIHe190108, FIHe190123 - FIHe190129 Observed with clear skies and sub-arcsec seeing. Observed
69-502 2024-06-09 01:23 Mon 24sex, hd99706, hd131496, hd108863,
hd192699, hd200964, hd4313
FIHf090109 - FIHf090117 ; FIHf090133 - FIHf090141 Thick clouds covering the sky at the beginning of the night that made it unable to observe any target until around 23:30 UT. HD99706 was observed twice; the first try failed to meet the requirements but then the clouds turned thinner and we could observe it according to the requirements. Managed to observe all targets except 24sex (missed due to the clouds). The rest of the targets were observed with some thin clouds and seeing of around 1 arcsec. Observed


The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: March 11 2024